In an arcade video game, a spot is programmed to move across the screen according to x = 6.29t - 0.683t3, where x is the distance in centimeters measured from the left edge of the screen and t is time in seconds. When the spot reaches a screen edge, either at x = 0 or x = 8.80 cm, t is reset to 0 and the spot starts moving again according to x(t). (a) At what time after starting is the spot instantaneously at rest? (b) At what value of x does this occur? (c) What is the spot's acceleration when this occurs? (d) At what time t > 0 does the spot first reach an edge of the screen?
Answer #1

Given that,

x = 6.29t - 0.683t3

(a)

v = dx / dt = d(6.29t - 0.683t3) / dt

v = 6.29 - 2.049t2

0 = 6.29 - 2.049t2

t = 1.752 s

(b)

Value of x,

x = 6.29t - 0.683t3

x = 6.29*1.752 - 0.683 * (1.752)3

x = 7.34 cm

(c)

spot's acceleration,

a = dv / dt = d( 6.29 - 2.049t2) / dt

a = -4.098t

a = -4.098 * 1.752

a = -7.179 cm/s^2

(d)

When the spot reach an edge of the screen,

x = 6.29t - 0.683t3

0 = t (6.29 - 0.683t2)

6.29 - 0.683t2 = 0

t = 3.034 s